OpenStudy (anonymous):

Solve for x log(x+2)+log(x-4)=log55

OpenStudy (therealmeeeee):

2 log (55/26) + log (28/175) = 2(log 5 + log 11 - log 2 - log 13) + log 4 + log 7 - log 7 - log 25 = 2 log 5 + 2 log 11 - 2 log 2 - 2 log 13 + 2 log 2 - 2 log 5 = 2 log 11 - 2 log 13 = 2x - 2y

OpenStudy (anonymous):

don't listen to the idiot above..

OpenStudy (anonymous):

Here's a hint: log(x+2)+log(x+4) = log((x+2)(x+4))

OpenStudy (therealmeeeee):

hold up rewind

OpenStudy (therealmeeeee):

dont call me no idiot lil ugley freak

OpenStudy (therealmeeeee):

we alll know whos smarter than yo lil dum self

OpenStudy (therealmeeeee):

hahaha u got 0 fans lame punk

OpenStudy (anonymous):

No need t argue I'm sure your both pretty smart.

OpenStudy (anonymous):

I have u solve for x and I beef on this problem forever could someone please help me out.

OpenStudy (anonymous):

@ECE you are right dont listen to @TheRealMeeeee He copied it from http://answers.yahoo.com/question/index?qid=20090301062638AASQMKR

OpenStudy (therealmeeeee):

no i did not